OpenStudy (anonymous):

The height of a parallelogram is 7 units more than the base. The measure of the base of the parallelogram is 7 units. What is the area, in square units, of the parallelogram? 343 49 98 171

OpenStudy (anonymous):

I think it's 98. I think I'm wrong though.

OpenStudy (ivettef365):

h = b + 7 b = 7 so h= 7+ 7 = 14 A = b.h

OpenStudy (ivettef365):

so you are right 14 x 7 = 98

OpenStudy (anonymous):

Yaaay! :D Thanks Ivette.

OpenStudy (ivettef365):

yw :)